\begin{thebibliography}{BDPS17}
	\expandafter\ifx\csname url\endcsname\relax
	\def\url#1{\texttt{#1}}\fi
	\expandafter\ifx\csname doi\endcsname\relax
	\def\doi#1{\burlalt{doi:#1}{http://doi.org/#1}}\fi
	\expandafter\ifx\csname urlprefix\endcsname\relax\def\urlprefix{URL }\fi
	\expandafter\ifx\csname href\endcsname\relax
	\def\href#1#2{#2}\fi
	\expandafter\ifx\csname burlalt\endcsname\relax
	\def\burlalt#1#2{\href{#2}{#1}}\fi
	
	\bibitem[AAB02]{HN-reduction-Anchouche-Azad-Biswas-2002}
	Boudjemaa Anchouche, Hassan Azad, and Indranil Biswas.
	\newblock Harder-{N}arasimhan reduction for principal bundles over a compact
	{K}\"ahler manifold.
	\newblock {\em Math. Ann.}, 323(4):693--712, 2002.
	\newblock \doi{10.1007/s002080200322}.
	
	\bibitem[AB01]{Anchouche-Biswas-2001}
	Boudjemaa Anchouche and Indranil Biswas.
	\newblock Einstein-{H}ermitian connections on polystable principal bundles over
	a compact {K}\"ahler manifold.
	\newblock {\em Amer. J. Math.}, 123(2):207--228, 2001.
	\newblock
	\urlprefix\url{http://muse.jhu.edu/journals/american_journal_of_mathematics/v123/123.2anchouche.pdf}.
	
	\bibitem[AB02]{Azad-Biswas-2002}
	Hassan Azad and Indranil Biswas.
	\newblock On holomorphic principal bundles over a compact {R}iemann surface
	admitting a flat connection.
	\newblock {\em Math. Ann.}, 322(2):333--346, 2002.
	\newblock \doi{10.1007/s002080100273}.
	
	\bibitem[AO24]{Alfaya-Oliveire-2024}
	David Alfaya and Andr\'{e} Oliveira.
	\newblock Lie algebroid connections, twisted {H}iggs bundles and motives of
	moduli spaces.
	\newblock {\em J. Geom. Phys.}, 201:Paper No. 105195, 55, 2024.
	\newblock \doi{10.1016/j.geomphys.2024.105195}.
	
	\bibitem[Ati57]{Atiyah-1957}
	M.~F. Atiyah.
	\newblock Complex analytic connections in fibre bundles.
	\newblock {\em Trans. Amer. Math. Soc.}, 85:181--207, 1957.
	\newblock \doi{10.2307/1992969}.
	
%	\bibitem[BBN05]{Balaji-Biswas-Nagaraj-2005}
%	V.~Balaji, Indranil Biswas, and D.~S. Nagaraj.
%	\newblock Krull-{S}chmidt reduction for principal bundles.
%	\newblock {\em J. Reine Angew. Math.}, 578:225--234, 2005.
%	\newblock \doi{10.1515/crll.2005.2005.578.225}.
	
	\bibitem[BDP18]{Biswas-Dan-Paul-2018}
	Indranil Biswas, Ananyo Dan, and Arjun Paul.
	\newblock Criterion for logarithmic connections with prescribed residues.
	\newblock {\em Manuscripta Math.}, 155(1-2):77--88, 2018.
	\newblock \doi{10.1007/s00229-017-0935-6}.
	
	\bibitem[BDPS17]{Biswas-Dan-Paul-Saha-2017}
	Indranil Biswas, Ananyo Dan, Arjun Paul, and Arideep Saha.
	\newblock Logarithmic connections on principal bundles over a {R}iemann
	surface.
	\newblock {\em Internat. J. Math.}, 28(12):1750088, 18, 2017.
	\newblock \doi{10.1142/S0129167X17500884}.
	
	\bibitem[BKS24]{Biswas-Kumar-Singh-2024}
	Indranil Biswas, Pradip Kumar, and Anoop Singh.
	\newblock A criterion for {L}ie algebroid connections on a compact {R}iemann
	surface.
	\newblock {\em Geom. Dedicata}, 218(4):Paper No. 87, 2024.
	\newblock \doi{10.1007/s10711-024-00938-8}.
	
	\bibitem[BP17]{Biswas-Paul-2017}
	Indranil Biswas and Arjun Paul.
	\newblock Equivariant bundles and connections.
	\newblock {\em Ann. Global Anal. Geom.}, 51(4):347--358, 2017.
	\newblock \doi{10.1007/s10455-016-9538-9}.
	
	\bibitem[GP20]{Gurjar-Paul-2020}
	Sudarshan Gurjar and Arjun Paul.
	\newblock Criterion for existence of a logarithmic connection on a principal
	bundle over a smooth complex projective variety.
	\newblock {\em Ann. Global Anal. Geom.}, 58(3):241--251, 2020.
	\newblock \doi{10.1007/s10455-020-09723-8}.
	
	\bibitem[HL10]{Huybrechts-Lehn-2010}
	Daniel Huybrechts and Manfred Lehn.
	\newblock {\em The geometry of moduli spaces of sheaves}.
	\newblock Cambridge Mathematical Library. Cambridge University Press,
	Cambridge, second edition, 2010.
	\newblock \doi{10.1017/CBO9780511711985}.
	
	\bibitem[Ram75]{Ramanathan-1975}
	A.~Ramanathan.
	\newblock Stable principal bundles on a compact {R}iemann surface.
	\newblock {\em Math. Ann.}, 213:129--152, 1975.
	\newblock \doi{10.1007/BF01343949}.
	
	\bibitem[Ram96]{Ramanathan-I}
	A.~Ramanathan.
	\newblock Moduli for principal bundles over algebraic curves. {I}.
	\newblock {\em Proc. Indian Acad. Sci. Math. Sci.}, 106(3):301--328, 1996.
	\newblock \doi{10.1007/BF02867438}.
	
	\bibitem[Wei38]{Weil-1938}
	Andre~A. Weil.
	\newblock Généralisation des fonctions abéliennes.
	\newblock {\em J. Math. Pures Appl.}, 17:47--87, 1938.
	
\end{thebibliography}

\documentclass[12pt,reqno]{amsart}
\usepackage[skip=5pt plus1pt, indent=20pt]{parskip}

\usepackage{palatino}
\usepackage{mathpazo}
\usepackage{amsmath,amssymb,amsxtra,color,calligra,mathrsfs,tcolorbox,stmaryrd}

\usepackage{xcolor} 
\colorlet{mdtRed}{red!50!black}
\colorlet{dblue}{blue!50!black}
\usepackage[colorlinks,pagebackref=true]{hyperref}
\hypersetup{linkcolor=dblue,citecolor=dblue,filecolor=dullmagenta,urlcolor=mdtRed}
\renewcommand*{\backref}[1]{}
\renewcommand*{\backrefalt}[4]{[{%
		\ifcase #1 Not cited.%
		\or $\uparrow$~#2.%
		\else $\uparrow$~#2.%
		\fi%
	}]}
\usepackage[all]{xy}
\usepackage{tikz,tikz-cd,tkz-graph,enumerate}
\usetikzlibrary{matrix,arrows,decorations.pathmorphing}

\usepackage[us,12hr]{datetime}
\usepackage{comment}

\DeclareMathOperator{\Id}{{\rm Id}}
\DeclareMathOperator{\Hom}{{\rm Hom}}

\DeclareMathOperator{\End}{{\rm End}}

\DeclareMathOperator{\Lie}{{\rm Lie}}
\DeclareMathOperator{\At}{{\rm At}}

\DeclareMathOperator{\ad}{{\rm ad}}

\DeclareMathOperator{\GL}{{\rm GL}}

\newcommand{\mf}[1]{\mathfrak{#1}}
\newcommand{\mc}[1]{\mathcal{#1}}

\newcommand{\bb}[1]{\mathbb{#1}}

\renewcommand{\ker}{\mathrm{Ker}}

\newcommand{\doi}[1]{\href{https://doi.org/#1}{doi:#1}}
\numberwithin{equation}{subsection}

\newtheorem{theorem}[equation]{Theorem}

\newtheorem{proposition}[equation]{Proposition}

\theoremstyle{definition}
\newtheorem{definition}[equation]{Definition}

\theoremstyle{theorem}

\makeatletter
\newcommand\fnsymb[1]{\textsuperscript{\@fnsymbol{#1}}}
\newcommand\fnletter[1]{\lowercase{\textsuperscript{\@alph{#1}}}}

\makeatother

\usepackage{marvosym} 
\renewcommand{\email}[2][1]{\thanks{\textit{Email address}#1: \href{mailto:#2}{#2}}}

\renewcommand{\address}[2][1]{\thanks{\textit{Address}#1: #2}} 

\begin{document}

\baselineskip=15.5pt 

\title[On Lie algebroid connections]{Lie Algebroid Connections on Principal Bundles}

\author[S. Ghosh]{Samit Ghosh\fnsymb{1}}

\author[A. Paul]{Arjun Paul\fnsymb{2}}

\email[\fnsymb{1}]{sg23rs005@iiserkol.ac.in}

\email[\fnsymb{2}]{arjun.paul@iiserkol.ac.in}

\address[]{Department of Mathematics and Statistics, 
	Indian Institute of Science Education and Research Kolkata, 
	Mohanpur - 741 246, Nadia, West Bengal, India.
}

\thanks{Corresponding author: Arjun Paul}

\subjclass[2010]{14J60, 53C07, 32L10}

\keywords{Connection; Lie algebroid; principal $G$--bundle.} 

\begin{abstract}
	Let $X$ be an irreducible smooth complex projective variety. 
	Let $G$ be a linear algebraic group over $\mathbb{C}$. 
	We define the notion of Lie algebroid valued connection on 
	holomorphic principal $G$--bundles on $X$, and study their 
	basic properties under extension and reduction of structure group. 
	Finally we investigate criterions for existence of a Lie algebroid 
	connection on principal $G$--bundles over smooth complex projective curves. 
\end{abstract}

\baselineskip=15.5pt 

\date{Last updated on \today\,at \currenttime\,(IST)}

\maketitle 

\tableofcontents

\section{Introduction}\label{sec:introduction}

A famous theorem of A. Weil \cite{Weil-1938} says that a holomorphic vector bundle $E$ 
on a compact connected Riemann surface $X$ admits a holomorphic connection if and only if 
each indecomposible holomorphic direct summand of $E$ has degree zero. 
In \cite{Atiyah-1957} M. Atiyah generalizes the notion of holomorphic connections 
in the context of holomorphic principal $G$--bundles on compact K\"ahler manifolds, 
and gives an algebro-geometric proof and Weils' theorem for holomorphic vector bundles 
on compact connected Riemann surfaces. 
In \cite{Azad-Biswas-2002} Azad and Biswas generalize Weils theorem for 
holomorphic principal $G$--bundles on compact connected Riemann surfaces. 
It is clear from these results that not every holomorphic vector bundles 
and principal $G$--bundles can admit holomorphic connections. 
This naturally leads one to consider the notion of meromorphic connections. 
One of the simplest kind of meromorphic connections is the notion of 
logarithmic connection, which are treated for holomorphic vector bundles 
and holomorphic principal $G$--bundles, for example in 
\cite{Biswas-Dan-Paul-2018}, \cite{Biswas-Dan-Paul-Saha-2017}, \cite{Gurjar-Paul-2020} etc. 

In the context of complex algebraic and differential geometry, the classical notion 
of holomorphic as-well-as singular connections has natural generalization by 
replacing tangent bundle with a Lie algebroid leading to the notion of 
{\it Lie algebroid connections}, which is more convenient to work in some setups 
like Poisson geometry, foliation theory etc. The notion of Lie algebroid connections 
also generalize the notion of holomorphic and logarithmic connections. 
It is an interesting problem to study Lie algebroid connections on holomorphic 
vector bundles and principal bundles. 

Let $X$ be a connected compact Riemann surface. Fix a holomorphic Lie algebroid 
$\mc V = (V, [\cdot\,,\,\cdot], \varphi)$ on $X$ with $V$ a stable vector bundle. 
In \cite{Biswas-Kumar-Singh-2024} the authors shows that every holomorphic vector bundle 
on $X$ admits a $\mc V$--valued Lie algebroid connection generalizing a result 
\cite[Corollary 3.17]{Alfaya-Oliveire-2024} of Alfaya and Oliveire. 
In this paper we generalize the notion of $\mc V$--valued Lie algebroid connections 
in the context of principal $G$--bundles 
(see Definition \ref{def:Lie-alg-conn-on-G-Bundles}), 
study their properties under extension and reduction of the structure group 
of the principal bundles (see \S\,\ref{sec:basic-properties}), 
and prove the following. 

\begin{theorem}
	Let $X$ be an irreducible smooth complex projective curve of genus $g \geq 2$. 
	Fix a Lie algebroid $\mc V = (V, [\cdot\,,\,\cdot], \varphi)$ on $X$ 
	such that $V$ is a stable vector bundle on $X$ with the slope $\mu(V) \neq 2-2g$. 
	Let $G$ be a reductive linear algebraic group over $\bb C$. 
	Then any holomorphic principal $G$--bundle $E_G$ on $X$ admits a 
	$\mc{V}$--valued Lie algebroid connection. 
\end{theorem}

This generalize the main result of \cite{Biswas-Kumar-Singh-2024} to 
the case of holomorphic principal $G$--bundles on $X$.

\section{Lie Algebroid Connections}

\subsection{The case of vector bundles}
Let $X$ be an irreducible smooth projective variety over $\mathbb{C}$. 
Let $\mc O_X$ be the sheaf of holomorphic functions on $X$, and let 
$TX$ be the holomorphic tangent bundle of $X$. 

\begin{definition}\cite[\S\,1.1]{Alfaya-Oliveire-2024}
	A {\it Lie algebroid} on $X$ is a triple $\mc V := (V, [\cdot, \cdot], \varphi)$, where 
	\begin{enumerate}[(i)]
		\item $V$ is a holomorphic vector bundle on $X$, 
		
		\item $[\cdot, \cdot] : V\times V \to V$ is a $\bb C$--bilinear skew-symmetric 
		morphism of sheaves such that for all locally defined sections 
		$u, v, w$ of $V$, the following Jacobi identity holds: 
		$$[u, [v, w]]+[v, [w, u]]+[w, [u, v]] = 0;$$
		
		\item $\varphi : V \to TX$ is a vector bundle homomorphism 
		satisfying the following properties: for all locally defined sections $s, t$ of $V$ and 
		locally defined section $f$ of $\mc O_X$, we have 
		\begin{enumerate}[(a)]
			\item {\it Compatibility of Lie algebra structures}: 
			$\varphi([s, t]) = [\varphi(s), \varphi(t)]$, and 
			\item {\it Leibniz rule}: $[fs, t] = f[s, t] - \varphi(t)(f)s$.
		\end{enumerate}
	\end{enumerate}
	The homomorphism $\varphi$ is called the {\it anchor map} of the Lie algebroid $\mc V$. 
	The {\it degree} and the {\it rank} of $\mc V$ is defined to be the degree and the rank, 
	respectively, of the underlying vector bundle $V$ of $\mc V$. 
\end{definition}

The dual of the anchor map gives a holomorphic vector bundle homomorphism 
$$\varphi^* : \Omega_X^1 \longrightarrow V^*,$$
where $\Omega_X^1$ is the holomorphic cotangent bundle of $X$. 
Fix a Lie algebroid $\mc V := (V, [\cdot, \cdot], \varphi)$ on $X$. 
Let $\mc E$ be a holomorphic vector bundle on $X$. 

\begin{definition}\label{def:Lie-alg-conn-on-VB}
	A {\it $\mc V$--valued Lie algebroid connection} on  
	$\mc E$ on $X$ is a $\bb C$--linear homomorphism of sheaves 
	$$D : \mc E \longrightarrow \mc E\otimes V^*$$ 
	satisfying the {\it $\varphi^*$--twisted Leibniz rule}: 
	\begin{equation}\label{eqn:phi^*-twisted-Leibniz-rule}
		D(f\cdot s) = fD(s)+s\otimes\varphi^*(df), 
	\end{equation}
	for all locally defined section $s$ of $\mc E$ and for all locally defined 
	section $f$ of $\mc O_X$. 
\end{definition}

\subsection{The case of principal $G$--bundles}
Now we extend the definition of Lie algebroid connection to the case of principal bundles 
following a construction given in \cite{Biswas-Paul-2017}. 
Let $G$ be a linear algebraic group over $\mathbb{C}$ with the Lie algebra $\mf{g} := \Lie(G)$. 
Let $p : E_G \to X$ be a holomorphic principal $G$--bundle on $X$. The adjoint representation 
$${\rm ad} : G \longrightarrow \GL(\mf{g})$$ 
of $G$ on its Lie algebra $\mf{g}$ gives rise to a vector bundle 
$$\ad(E_G) := E_G\times^{\rm ad}\mf{g}$$ 
on $X$, called the {\it adjoint vector bundle} of $E_G$. 
If $E$ is the frame bundle of a vector bundle $\mc E$ of rank $n$ on $X$, 
then we have $\ad(E) \cong \End(\mc E)$, the endomorphism bundle of $\mc E$. 
The surjective submersion $p : E_G \to X$ gives rise to an exact sequence of 
vector bundles 
\begin{equation}
	\xymatrix{
		0 \ar[r] & \ad(E_G) \ar[r] & \At(E_G) \ar[r]^-{d'p} & TX \ar[r] & 0 
	}
\end{equation}
called the Atiyah exact sequence of $E_G$. 
A connection on the principal $G$--bundle $E_G$ on $X$ is an $\mc O_X$--linear 
homomorphism $\nabla : TX \to \At(E_G)$ such that $d'p\circ\nabla = \Id_{TX}$.

Fix a Lie algebroid $\mc V = (V, [\cdot\,,\,\cdot], \varphi)$ on $X$, 
and consider the map 
$$\rho : \At(E_G)\oplus V \longrightarrow TX$$ 
defined by 
\begin{equation}
	\rho(\xi, v) = d'p(\xi)-\varphi(v), 
\end{equation}
for all locally defined section $\xi$ of $\At(E_G)$ and locally defined section 
$v$ of $V$. Note that $\rho$ is a vector bundle homomorphism and 
\begin{equation}
	\At_{\varphi}(E_G) := \rho^{-1}(0) 
\end{equation}
is a vector bundle on $X$. 
The restriction of the second projection map gives rise 
to a vector bundle homomorphism 
\begin{equation}\label{eqn:rho-tilde-map}
	\widetilde{\rho} : \At_{\varphi}(E_G) \longrightarrow V
\end{equation} 
with kernel 
$$\ker(\widetilde{\rho}) = \ad(E_G).$$ 
Thus we have the following short exact sequence 
\begin{equation}\label{eqn:Atiyah-exact-seq-for-(V,phi)-valued-connection}
	0 \longrightarrow \ad(E_G) \longrightarrow \At_{\varphi}(E_G) 
	\stackrel{\widetilde{\rho}}{\longrightarrow} V \longrightarrow 0 
\end{equation}
of vector bundles on $X$, which fits into the following 
commutative diagram 
\begin{equation}
	\begin{gathered}
		\xymatrix{
		0 \ar[r] & \ad(E_G) \ar[r] \ar@{=}[d] & \At_{\varphi}(E_G) \ar[r]^-{\widetilde{\rho}} \ar[d] 
		& V \ar[r] \ar[d]^-\varphi & 0 \\ 
		0 \ar[r] & \ad(E_G) \ar[r] & \At(E_G) \ar[r]^-{d'p} & TX \ar[r] & 0 
		}
	\end{gathered}
\end{equation}
of vector bundle homomorphisms with all rows exact. 

\begin{definition}\label{def:Lie-alg-conn-on-G-Bundles}
	A {\it $\mc V$--valued Lie algebroid connection} on $E_G$ is a 
	vector bundle homomorphism 
	$$\nabla : V \longrightarrow \At_{\varphi}(E_G)$$
	such that $\widetilde{\rho}\circ\nabla = \Id_V$, 
	where $\widetilde{\rho}$ is defined in \eqref{eqn:rho-tilde-map}. 
\end{definition}

The short exact sequence \eqref{eqn:Atiyah-exact-seq-for-(V,phi)-valued-connection} 
defines a cohomology class 
\begin{equation}\label{defn:V-valued-Atiyah-class-of-E_G}
	\Phi_{\mc V}(E_G) \in H^1(X, \ad(E_G)\otimes V^*), 
\end{equation}
such that the exact sequence \eqref{eqn:Atiyah-exact-seq-for-(V,phi)-valued-connection} 
splits holomorphically if and only if $\Phi_{\mc V}(E_G) = 0$. 

\begin{proposition}\label{prop:E_G-admits-Lie-alg-conn-iff-Atiyah-cls-vanishes}
	A holomorphic principal $G$--bundle $E_G$ on $X$ admits a $\mc V$--valued 
	holomorphic Lie algebroid connection if and only if $\Phi_{\mc V}(E_G) = 0$. 
	We call $\Phi_{\mc V}(E_G)$ the {\it $\mc V$--valued Atiyah class of $E_G$}. 
\end{proposition}

Let $\nabla : V \to \At_{\varphi}(E_G)$ be a $\mc V$--valued Lie algebroid connection 
on $E_G$ over $X$. For all locally defined holomorphic sections $s$ and $t$ of $V$, let 
$$\kappa_{\nabla}(s, t) := [\nabla(s), \nabla(t)] - \nabla([s, t]).$$ 
Since the homomorphism $\widetilde{\rho} : \At_{\varphi}(E_G) \to V$ respects the Lie algebra 
structures on the sheaves of sections, $\kappa_{\nabla}(s, t)$ defines a holomorphic local 
section of $\ad(E_G)$. Thus we obtain a section 
$$\kappa_{\nabla} \in H^0(X, \ad(E_G)\otimes \bigwedge\nolimits^2V^*),$$
called the {\it curvature} of the $\mc V$--valued Lie algebroid connection 
$\nabla$ on $E_G$. The section $\kappa_{\nabla}$ can be considered as an obstruction 
for $\nabla$ to be a Lie algebra homomorphism. 

\begin{definition}
	A $\mc V$--valued Lie algebroid connection $\nabla$ on a 
	principal $G$--bundle $E_G$ on $X$ is said to be {\it flat} if $\kappa_{\nabla} = 0$. 
\end{definition}

\begin{proposition}
	If ${\rm rank}(\mc V) = 1$, any $\mc V$--valued 
	Lie algebroid connection on $E_G$ is flat. 
\end{proposition}

\begin{proof}
	If ${\rm rank}(\mc V) = 1$, then $\bigwedge\nolimits^2 V^* = 0$ and so for any 
	$\mc V$--valued connection $\nabla$ on $E_G$, its curvature 
	$\kappa_{\nabla}$, being an element of $H^0(X, \ad(E_G)\otimes \bigwedge\nolimits^2V^*) = 0$, 
	vanishes identically. This completes the proof. 
\end{proof}

\section{Basic Properties}\label{sec:basic-properties}

\subsection{Extension of structure groups}
Let $G$ and $H$ be linear algebraic groups over $\bb C$ with their Lie algebras 
$\mf g$ and $\mf h$, respectively. Given a homomorphism of algebraic groups $f : G \to H$ 
over $\bb C$, let $df : \mf g \to \mf h$ be the Lie algebra homomorphism induced by $f$. 
Let $p : E_G \to X$ be a holomorphic principal $G$--bundle over $X$, and let 
$$p' : E_H := E_G\times^f H \to X$$ 
be the associated principal $H$--bundle on $X$ obtained by extending the structure group 
of $E_G$ along $f$. Let 
\begin{align*}
	& \ad(f) : \ad(E_G) \longrightarrow \ad(E_H) \\ 
	\text{and} \ \ 
	& \At(f) : \At(E_G) \longrightarrow \At(E_H) 
\end{align*}
be the homomorphisms of the adjoint bundles and the Atiyah bundles of $E_G$ and $E_H$, 
respectively, induced by $f$. Then we have the following commutative diagram of vector bundle 
homomorphisms 
\begin{equation}\label{diag:comm-diag-of-Atiyah-exact-seqn-for-extn-of-str-grp}
	\begin{gathered}
		\xymatrix{
			0 \ar[r] & \ad(E_G) \ar[d]^{\ad(f)} \ar[r]^{\iota_G} & \At(E_G) \ar[d]^{\At(f)} \ar[r]^-{d'p} 
			& TX \ar@{=}[d] \ar[r] & 0 \\ 
			0 \ar[r] & \ad(E_H) \ar[r]^{\iota_H} & \At(E_H) \ar[r]^-{d'p'} & TX \ar[r] & 0\,. 
		}
	\end{gathered}
\end{equation}
It is clear from the above diagram that a holomorphic connection on $E_G$ induces a 
holomorphic connection on $E_H := E_G\times^fH$. 
Let $$\rho' : \At(E_H)\oplus V \to TX$$ be the homomorphism defined by 
$$\rho'(\xi, v) = d'p'(\xi) - \varphi(v),$$
for all locally defined sections $\xi$ of $\At(E_H)$ and $v$ of $V$, respectively. 
Let $$\widetilde{\rho'} : \At_{\varphi}(E_H) := \ker(\rho') \longrightarrow V$$ 
be the restriction of the second projection map. 
Then we have a vector bundle homomorphism 
$$\At_{\varphi}(f) : \At_{\varphi}(E_G) \longrightarrow \At_{\varphi}(E_H)$$ 
such that $\widetilde{\rho'}\circ\At_{\varphi}(f) = \widetilde{\rho}$. 
Thus, the above commutative diagram 
\eqref{diag:comm-diag-of-Atiyah-exact-seqn-for-extn-of-str-grp} induces the 
following commutative diagram of vector bundles and homomorphisms 
\begin{equation}\label{diag:phi-twisted-diag-of-Atiyah-exact-seqns-for-extn-of-str-grp}
	\begin{gathered}
		\xymatrix{
			0 \ar[r] & \ad(E_G) \ar[d]^{\ad(f)} \ar[r] & \At_{\varphi}(E_G) 
			\ar[d]^{\At_{\varphi}(f)} \ar[r]^-{\widetilde{\rho}} & V \ar@{=}[d] \ar[r] & 0 \\ 
			0 \ar[r] & \ad(E_H) \ar[r] & \At_{\varphi}(E_H) \ar[r]^-{\widetilde{\rho'}} & 
			V \ar[r] & 0\,. 
		}
	\end{gathered}
\end{equation}
From this, we have a natural homomorphism of cohomologies 
\begin{equation}
	H^1(f) : H^1(X, \ad(E_G)\otimes V^*) \longrightarrow H^1(X, \ad(E_H)\otimes V^*) 
\end{equation}
such that $H^1(f)(\Phi_{\mc V}(E_G)) = \Phi_{\mc V}(E_H)$. 
As an immediate consequence of it, we have the following result.

\begin{proposition}\label{prop:Lie-algebroid-connection-under-extn-of-str-grp}
	Let $f : G \to H$ be a homomorphism of linear algebraic groups over $\mathbb{C}$. 
	Let $E_G$ be a holomorphic principal $G$--bundle on $X$, and let $E_H$ be the 
	holomorphic principal $H$--bundle on $X$ obtained from $E_G$ by extension of its 
	structure group along $f$. Then any $\mc V$--valued Lie algebroid connection on 
	$E_G$ induces a $\mc V$--valued Lie algebroid connection on $E_H$.
\end{proposition}

\begin{proof}
	If $E_G$ admits a $\mc V$--valued Lie algebroid connection, 
	then $\Phi_{\mc V}(E_G) = 0$. 
	Since $\Phi_{\mc V}(E_H) = H^1(f)(\Phi_{\mc V}(E_G)) = 0$, 
	the result follows from 
	Proposition \ref{prop:E_G-admits-Lie-alg-conn-iff-Atiyah-cls-vanishes}
\end{proof}

\subsection{Reduction of structure group}

Now it is interesting to ask the following question: 
Suppose that $f : G \to H$ be a homomorphism of linear algebraic groups over $\bb C$. 
If $E_H$ admits a $\mc V$--valued connection, does $E_G$ admits a $\mc V$--valued connection? 
We give partial answers to this question. 

\begin{proposition}\label{prop:Key-reduction-for-reductive-subgroup}
	Let $f : G \to H$ be an injective homomorphism of linear algebraic groups over $\bb C$ 
	with $G$ reductive. Let $E_G$ be a principal $G$--bundle on $X$, and let 
	$$E_H = E_G\times^f H$$ 
	be the principal $H$--bundle on $X$ obtained by extending the structure group of 
	$E_G$ along the homomorphism $f$. 
	If $E_H$ admits a $\mc V$--valued Lie algebroid connection, then $E_G$ admits a 
	$\mc V$--valued Lie algebroid connection. 
\end{proposition}

\begin{proof}
	Let $\mf g := {\rm Lie}(G)$ and $\mf h := {\rm Lie}(H)$ be the Lie algebras of $G$ 
	and $H$, respectively. Let $df : \mf g \to \mf h$ be the Lie algebra homomorphism 
	induced by $f$, and let 
	$$\ad(f) : \ad(E_G) \longrightarrow \ad(E_H)$$
	be the vector bundle homomorphism induced by $df$.  
	Let $\alpha : G \to \End(\mf g)$ and $\beta : H \to \End(\mf h)$ be the 
	adjoint actions of $G$ and $H$, respectively, on their Lie algebras. 
	Then the composite map 
	$$\beta\circ f : G \to \End(\mf h)$$ 
	gives an adjoint action of $G$ on $\mf h$. 
	Since $df$ is a $G$--module homomorphism and $G$ is reductive, 
	there is a $G$--submodule $W$ of $\mf h$ such that 
	\begin{equation}\label{eqn:direct-sum-decomp}
		\mf h = df(\mf g)\bigoplus W 
	\end{equation}
	as $G$--modules. Since $df$ is injective, from the direct sum decomposition of 
	$G$--modules in \eqref{eqn:direct-sum-decomp} projecting to the first factor 
	we get a $G$--module homomorphism $\pi_{\mf g} : \mf h \to \mf g$ such that 
	$\pi_{\mf g}\circ df = \Id_{\mf g}$. Then $\pi_{\mf g}$ induces a vector bundle 
	homomorphism 
	\begin{equation}
		\widetilde{\pi_{\mf g}} : \ad(E_H) \longrightarrow \ad(E_G) 
	\end{equation}
	such that $\widetilde{\pi_{\mf g}}\circ\ad(f) = \Id_{{\rm ad}(E_G)}$. 
	
	Suppose that $E_H$ admits a $\mc V$--valued Lie algebroid connection. 
	Then there exists a $\mc O_X$--module homomorphism 
	$$\eta : \At_{\varphi}(E_H) \to \ad(E_H)$$ 
	such that $\eta\circ\iota_H = \Id_{\ad(E_H)}$, 
	where $\iota_H : \ad(E_H) \to \At_{\varphi}(E_H)$ is the homomorphism 
	in \eqref{diag:comm-diag-of-Atiyah-exact-seqn-for-extn-of-str-grp}. 
	Then it follows from the commutativity of the diagram 
	\eqref{diag:comm-diag-of-Atiyah-exact-seqn-for-extn-of-str-grp} 
	that the composition 
	$$\widetilde{\pi_{\mf g}}\circ\eta\circ\At(f) : \At_{\varphi}(E_G) \to \ad(E_G)$$
	gives an $\mc O_X$--linear splitting of the top exact sequence in 
	\eqref{diag:comm-diag-of-Atiyah-exact-seqn-for-extn-of-str-grp}. 
	Thus $E_G$ admits a $\mc V$--valued Lie algebroid connection. 
\end{proof}

Now we consider the case when the structure group of a principal bundle is not reductive. 
Let $G$ be a reductive linear algebraic group over $\bb C$. 
A closed subgroup $P$ of $G$ is said to be {\it parabolic} if $G/P$ is a complete $\bb C$--variety. 
Let $P$ be a parbolic subgroup of $G$.  
Let $\mf{R}_u(P)$ be the unipotent radical of $P$, and let 
$$q : P \longrightarrow P/\mf{R}_u(P)$$ 
be the associated quotient map. Let $L \subseteq P$ be a {\it Levi factor of $P$}; 
a closed connected subgroup of $P$ such that $q\big\vert_L : L \to P/\mf{R}_u(P)$ 
is an isomorphism of algebraic groups over $\bb C$. Note that $L$ is reductive. 
Given a principal $P$--bundle $E_P$ on $X$, let $E_L := E_P\times^{q'}L$ be the 
principal $L$--bundle on $X$ obtained by extending the structure group of $E_P$ 
along the homomorphism $$q' := (q\big\vert_L)^{-1}\circ\,q : P \to L.$$ 
The action of $P$ on the nilpotent radical $\mf{n} := \Lie(\mf{R}_u(P))$ of the 
Lie algebra $\mf{p} := \Lie(P)$ gives rise to a subbundle 
$E_P(\mf{n}) := E_P\times^P\mf{n}$ of the adjoint bundle $\ad(E_P)$ of $E_P$, 
and the associated quotient bundle $\ad(E_P)/E_P(\mf{n}) \cong E_P(\mf{l}) = \ad(E_L)$, 
where $\mf{l} = \Lie(L)$ is the Lie algebra of $L$. 
Then we have the following commutative diagram of vector bundle homomorphisms with all 
rows and columns exact 
(c.f. \eqref{diag:phi-twisted-diag-of-Atiyah-exact-seqns-for-extn-of-str-grp}): 
\begin{equation}
	\begin{gathered}
		\xymatrix{
			& 0 \ar[d] & 0 \ar[d] & & & \\ 
			& E_P(\mf{n}) \ar@{=}[r] \ar[d] & E_P(\mf{n}) \ar[d] & & & \\ 
			0 \ar[r] & \ad(E_P) \ar[r] \ar[d]^-{\ad(q')} & \At_{\varphi}(E_P) \ar[d]^-{\At(q')} 
			\ar[r]^-{\widetilde{\rho}_P} & V \ar@{=}[d] \ar[r] & 0 \\ 
			0 \ar[r] & \ad(E_L) \ar[d] \ar[r] & \At_{\varphi}(E_L) \ar[d] 
			\ar[r]^-{\widetilde{\rho}_L} & V \ar[r] & 0 \\ 
			& 0 & 0 & &&& 
		}
	\end{gathered}
\end{equation}
Suppose that $E_L$ admits a $\mc V$--valued Lie algebroid connection 
$\nabla : V \to \At_{\varphi}(E_L)$. Then $\widetilde{\rho}_L\circ\nabla = \Id_V$. 
Then the subsheaf 
$$\mc{E}_\nabla := \At(q')^{-1}\left(\nabla(V)\right) \subseteq \At_{\varphi}(E_P)$$ 
fits into the following short exact sequence of $\mc{O}_X$--modules 
\begin{equation}\label{eqn:short-ext-seq-for-reduction-of-str-grp-connection}
	0 \to E_P(\mf{n}) \to \mc{E}_\nabla \to V \to 0 
\end{equation}
on $X$ whose splitting gives rise to a $\mc V$--valued connection on $E_P$. 
Note that the above short exact sequence 
\eqref{eqn:short-ext-seq-for-reduction-of-str-grp-connection} defines a cohomology class 
\begin{equation}
	\Phi(E_P, L, \nabla) \in H^1(X, E_P(\mf{n})\otimes V^*), 
\end{equation}
which vanishes if and only if the exact sequence in 
\eqref{eqn:short-ext-seq-for-reduction-of-str-grp-connection} splits $\mc{O}_X$--linearly. 
From this, we have the following result. 

\begin{proposition}\label{prop:connection-on-E_P-from-E_L}
	With the above notations, if $H^1(X, E_P(\mf{n})\otimes V^*) = 0$, then a 
	$\mc V$--valued Lie algebroid connection on $E_L$ gives rise to a 
	$\mc V$--valued Lie algebroid connection on $E_P$. 
\end{proposition}

\section{Existence of Lie Algebroid Connections}
In this section we assume that $X$ is an irreducible smooth complex projective curve 
of genus $g \geq 2$. The {\it degree} of a coherent sheaf of $\mc O_X$--modules $E$ on $X$ 
is defined by 
$$\deg(E) := \int_X c_1(E) \in \bb Z,$$
where $c_1(E)$ stands for the first Chern class of $E$. 
The rational number 
$$\mu(E) := \frac{\deg(E)}{{\rm rank}(E)}$$ 
is called the {\it slope} of $E$. 

\begin{definition}
	A vector bundle $E$ on $X$ is said to be {\it stable} (resp., {\it semistable}) 
	if for any non-zero proper subsheaf $F$ of $E$ we have 
	$\mu(F) < \mu(E)$ (resp., $\mu(F) \leq \mu(E)$). 
\end{definition}

The notion of slope semistablity and stability has a natural generalization to the case of 
principal $G$--bundles on $X$. Let $G$ be a reductive linear algebraic group over $\bb C$. 
If a principal $G$--bundle $E_G$ on $X$ admits a holomorphic reduction $E_P \subseteq E_G$ 
of its structure group to a parabolic subgroup $P \subseteq G$, for any character 
$\chi : P \to \bb G_m$ of $P$, we get a holomorphic line bundle 
$$\chi_*E_P := E_P\times^{\chi} \bb G_a$$ 
on $X$. 

\begin{definition}\cite{Ramanathan-I, Ramanathan-1975}
	A principal $G$--bundle $E_G$ on $X$ is said to be {\it semistable} (resp., {\it stable}) 
	if for any reduction $E_P \subseteq E_G$ of the structure group of $E_G$ to a proper 
	parabolic subgroup $P \subseteq G$, and any nontrivial dominant character 
	$\chi : P \to {\bb G}_m$, we have $\deg(\chi_*E_P) \leq 0 (\textnormal{resp., } < 0)$. 
\end{definition}

Fix a Lie algebroid $\mc V = (V, [\cdot, \cdot], \varphi)$ on $X$ 
such that the underlying vector bundle $V$ of $\mc V$ is stable. 
Let 
$$\mu(\mc V) := \frac{\deg(V)}{{\rm rank}(V)}$$ 
be the {\it slope} of the underlying vector bundle $V$ of the Lie algebroid $\mc V$. 
Note that $TX$ is a line bundle on $X$ with the slope $\mu(TX) = 2-2g$. 

If $\mu(\mc V) > 2-2g = \mu(TX)$, then both $V$ and $TX$ being stable vector bundles 
on $X$ we have $H^0(X, \Hom(V, TX)) = 0$ (see \cite[Proposition 1.2.7]{Huybrechts-Lehn-2010}), 
and hence $\varphi = 0$ in this case. Then a $\mc V$--valued Lie algebroid connection 
on $E_G$ is just a global section of $\ad(E_G)\otimes V^*$; so we may take the 
zero section in $H^0(X, \ad(E_G)\otimes V^*)$, in particular. 

If $\mu(\mc V) = 2-2g = \mu(TX)$, then any non-zero $\mc O_X$--module homomorphism 
$\varphi : V \to TX$ is an isomorphism (see \cite[Proposition 1.2.7]{Huybrechts-Lehn-2010}). 
Then we may replace $V$ with $TX$ so that a $\mc V$--valued Lie algebroid connection 
on $E_G$ is nothing but a holomorphic connection on $E_G$. This case is studied in 
detail in \cite{Azad-Biswas-2002}. 

Now we assume that $\mu(\mc V) < 2-2g = \mu(TX)$. Then we have the following. 

\begin{proposition}\label{prop:existence_semistable-G-bundles-on-curve}
	Let $G$ be a reductive linear algebraic group over $\bb C$. 
	With the above assumptions on $\mc V$, any semistable principal $G$--bundle 
	$E_G$ on $X$ admits a $\mc V$--valued Lie algebroid connection. 
\end{proposition}

\begin{proof}
	Let $\Phi_{\mc V}(E_G) \in H^1(X, \ad(E_G)\otimes V^*)$ be the $\mc V$--valued Atiyah class 
	of $E_G$. By Serre duality, we have 
	\begin{equation*}
		H^1(X, \ad(E_G)\otimes V^*) \cong H^0(X, \ad(E_G)^*\otimes V\otimes K_X)^*,
	\end{equation*} 
	where $K_X = \Omega_X^1$ is the canonical line bundle on $X$. 
	Since $E_G$ is semistable by assumption, its adjoint bundle $\ad(E_G)$ is semistable 
	by \cite[Proposition 2.10]{Anchouche-Biswas-2001}. 
	Then the tensor product bundle $\ad(E_G)^*\otimes V\otimes K_X$ is semistable 
	(see \cite[Theorem 3.1.4]{Huybrechts-Lehn-2010}). 
	Since $G$ is reductive, the adjoint bundle $\ad(E_G)$ is isomorphic to its dual, 
	and hence $\deg(\ad(E_G)) = 0$. Then we have 
	$$\mu(\ad(E_G)^*\otimes V\otimes K_X) = \mu(K_X)+\mu(V) = 2g-2+\mu(V)<0.$$
	Then by \cite[Proposition 1.2.7]{Huybrechts-Lehn-2010} $H^0(X, \ad(E_G)^*\otimes V\otimes K_X) = 0$, 
	and hence $\Phi_{\mc V}(E_G) = 0$. Hence the result follows. 
\end{proof}

\begin{theorem}\label{thm:main-thm}
	Fix a Lie algebroid $\mc V = (V, [\cdot\,,\,\cdot], \varphi)$ on $X$ 
	such that $V$ is stable with $\mu(V) < 2-2g = \deg(TX)$. 
	Let $G$ be a reductive linear algebraic group over $\bb C$. 
	Let $E_G$ be a principal $G$--bundle on $X$. 
	Then $E_G$ admits a $\mc V$--valued Lie algebroid connection. 
\end{theorem}

\begin{proof}
	Let $E_G$ be a principal $G$--bundle on $X$. Since $G$ is reductive, 
	by \cite[Theorem 1]{HN-reduction-Anchouche-Azad-Biswas-2002} 
	$E_G$ admits a canonical reduction 
	$E_P \subseteq E_G$ of its structure group to a parabolic subgroup 
	$P \subseteq G$ such that the associated principal $L$--bundle 
	$$E_L := E_P\times^q L$$ 
	obtained by extension of the structure group of $E_P$ by the 
	quotient homomorphism 
	$$q : P \longrightarrow P/{\mf R}_u(P) \cong L,$$ 
	is semistable; here $L$ is the {\it Levi factor} of $P$, 
	a closed connected reductive subgroup of $P$ such that 
	the restriction of the quotient homomorphism $q : P \to P/{\mf R}_u(P)$ 
	to $L \subseteq P$ is an isomorphism of algebraic groups over $\bb C$. 
	Then by Proposition \ref{prop:existence_semistable-G-bundles-on-curve} the principal 
	$L$--bundle $E_L$ admits a $\mc V$--valued Lie algebroid connection. 
	Since $\mu_{\min}(E_P(\mf n)) \geq 0$ 
	by \cite{HN-reduction-Anchouche-Azad-Biswas-2002} 
	and $V\otimes K_X$ is semistable with $\mu(V\otimes K_X) < 0$, 
	it follows that $\Hom(E_P(\mf n), V\otimes K_X) = 0$, 
	and hence $H^1(X, E_P(\mf n)\otimes V^*) = 0$ by Serre duality. 
	Then by Proposition \ref{prop:connection-on-E_P-from-E_L} 
	that $E_P$ admits a $\mc V$--valued Lie algebroid connection, and then 
	by Proposition \ref{prop:Lie-algebroid-connection-under-extn-of-str-grp} 
	$E_G$ admits a $\mc V$--valued Lie algebroid connection. 
	This completes the proof. 
\end{proof}

\section*{Acknowledgment}
The first named author is supported by the \textit{National Board of Higher Mathematics (NBHM)} 
through the Doctoral Research Fellowship Program. 
The second named author is partially supported by the DST INSPIRE Faculty Fellowship 
(Research Grant No.: DST/INSPIRE/04/2020/000649, IFA20-MA-144), the Ministry of Science \& Technology, 
Government of India. 


\end{document}